OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + y = 11 –3x + 2y = 1

jimthompson5910 (jim_thompson5910):

Hint: add the equations to get 3x + y = 11 –3x + 2y = 1 ------------ 0x + 3y = 12

OpenStudy (anonymous):

@junior_swaqq http://openstudy.com/users/junior_swaqq#/updates/51c0c5c1e4b0a49e8c8f7fc0 we just helped you in a similar one. please refer back to the steps in the previous one. then if you have questions, please ask :)